You perform the Simple Harmonic Motion experiment and create a plot of Period-Squared vs. Mass. You add a linear fit and find the following equation of the line: y=9.52x-0.067

What is the spring constant, k (in N/m or kg/s2)?
Answer #1

The time period of a SHM is

T = 2*pi * sqrt(m/k)

where m is the mass and k is the spring constant.

So, T^2 /m = 4*pi^2 /k

From the equation of the line, we can deduce that T^2 /m is nothing but the slope of the line.

So, T^2 /m = 9.52

So, 9.52 = 4*pi^2 /k

k = 4.15 N/m